Consider the reaction between calcium oxide and carbon dioxide.
CaO + CO2>>> CaCO3
A chemist allows 14.4g of CaO and 13.8g of CO2 to react. When the reaction is finished the chemist collects 19.4 of CaCO3. Determine the limiting reactant, theoritical yield and precent yield for the reaction.
(Molar masses, g/mol:CaO=56.08,CO2=28.01,CaCO3=100.09)
1
Expert's answer
2020-08-05T05:09:55-0400

from the number of moles of reactants and products, we can determine the % yield and thus establish the limiting reactant

number of moles is calculated from the ratio of mass and molar mass

number of moles of CaO = 14.4/56.1= 0.26 mol

nummber of moles of CO2= 13.8/44= 0.31 mol

to establish the limiting reactant, compare the number of moles of CaO, CO2 with that of CaCO3 produced :

the reacting mole ratios of reactants and products = 1:1:1

hence, the number of moles of CaCO3=0.26 mol

The number of moles of CaCO3 produced by CaO is less than the theoretical yield, hence CaO is the limiting reactant: if amount of CaO is increased then more of CaCO3 will be produced from the available CO2

mass of CaCO3 produced = 0.26*100=26 g where 100 = molar mass of CaCO3

% yield = (actual yield/theoretical yield)*100

= (19.4/26)*100

=74.6%